# nbi: the Astronomer’s Package for Neural Posterior Estimation

### 06 Dec 2023 v1

@@ 1 Problem Statement
The first impediment is the need for custom featurizer (a.k.a. embedding/compressor) networks tailored to the observed data, which compress the high-dimensional observed data into a low-dimensional vector to be “read in” by the NDE as the conditional. This customization need creates a steep learning curve for the domain expert less steeped in deep learning, which is nevertheless required under existing NPE software frameworks such as sbi (Tejero-Cantero et al., 2020) and LAMPE [^2] https://github.com/francois-rozet/lampe. However, this need for customization contrasts sharply with the limited data types that astronomers regularly face—namely sequential data like light curves and spectra, and spatial data (images), for which there exist generic featurizer network architectures that are proven to be effective.
@@ 2 The nbi Framework
Our framework is similar to existing frameworks in many ways. We support both ANPE and SNPE (for $n_{\rm rounds}>1$), and we utilize the Masked Autoregressive Flow (MAF; Papamakarios et al. 2017) as the NDE, whose hyper-parameters are specified via the dictionary flow during initialization. The MAF is adapted from our previous work (Zhang et al., 2021), which truncates the scaling factor for each MAF block to improve training stability. On the other hand, our framework differentiates from existing frameworks with the following features:
@@ 3 Asymptotically Exact Inference
The availability of tractable likelihood functions for vast astronomical inference problems contrasts sharply with problems traditionally tackled with simulation-based inference. At the core of nbi is the use of the likelihood function to re-weight the NDE surrogate posterior into asymptotically exact posteriors via importance sampling. The importance weight for each discrete sample of the NDE surrogate posterior is:
@@
SNPE: The integration of importance sampling with NPE is particularly interesting for SNPE. Under SNPE, the NDE surrogate posterior is trained over multiple rounds, where the surrogate posterior from the previous round is used to generate the training set of the next round. The key here is that the compute expended for simulating the next round training set can be directly used for calculating the importance weights for the current round NDE posterior. Thus, each round of SNPE contributes to the final effective sample size with no extra cost. The ratio between the effective and total sample size (simulation efficiency) for each round becomes an indicator for the convergence of the NDE posterior, and thus an early stopping criterion. The full SNPE with importance sampling integration (SNPE-IS) algorithm is presented in Algorithm 1.
@@
2:  Output: posterior samples $\{\theta_{1:NR}\}$, weights $\{w_{1:NR}\}$, effective sample sizes $\{n_{\rm eff,1:R}\}$, final surrogate posterior $\hat{q}(\theta)$
@@
14:     Train $\hat{q}(\theta)$ on $\{\theta_{1:N},x_{1:N}\}$ using NLL
@@ 4 Discussion
We consider the current version (v0.4.1) of nbi “production-ready” for inference problems involving sequential data, such as light curves and spectra. The only customization we require is a “noise” function that captures the instrumental effects of the underlying dataset. Looking forward, we plan to include more types of featurizer networks to allow for the out-of-the-box utility of nbi to be extended to more types of data. One aspect worth mentioning is specialized networks for irregularly sampled light curve. We point out that for single inference tasks utilizing SNPE, there is no distinction between regularly and irregularly sampled data, and the 1D ResNet-GRU network provided with nbi should be sufficient. We outline several aspects for an extended version of this work.

### 21 Dec 2023 v2

@@ 1 Problem Statement
The first impediment is the need for custom featurizer/embedding networks tailored to the observed data, which compress the high-dimensional observed data into a low-dimensional vector to be “read in” by the NDE as the conditional. This need for customization beyond standard convolutional and recurrent neural networks creates a steep learning curve for the domain expert less steeped in deep learning. On the other hand, there exist network architectures that are well benchmarked and demonstrated to be effective for common astronomical datasets (e.g. Becker et al. 2020; Zhang & Bloom 2021; Chan et al. 2021; Moreno-Cartagena et al. 2023, notably including sequential data like light curves and spectra.
@@ 2 The nbi Framework
Our framework is similar to existing frameworks such as sbi (Tejero-Cantero et al., 2020) and LAMPE [^3] https://github.com/francois-rozet/lampe in many ways. We support both ANPE and SNPE (for $n_{\rm rounds}>1$), and we utilize the Masked Autoregressive Flow (MAF; Papamakarios et al. 2017) as the NDE, whose hyper-parameters are specified via the dictionary flow during initialization. The MAF is adapted from our previous work (Zhang et al., 2021), which truncates the scaling factor for each MAF block to improve training stability. On the other hand, our framework differentiates from existing frameworks with the following features:
@@ 3 Asymptotically Exact Inference
The availability of tractable likelihood functions for vast astronomical inference problems contrasts sharply with problems traditionally tackled with simulation-based inference. At the core of nbi is the use of the likelihood function to re-weight the NDE surrogate posterior into asymptotically exact posteriors via importance sampling (Paige & Wood, 2016; Müller et al., 2019; Dax et al., 2023). The importance weight for each discrete sample of the NDE surrogate posterior is:
@@
SNPE: The integration of importance sampling with NPE is particularly interesting for SNPE. Under SNPE, the NDE surrogate posterior is trained over multiple rounds, where the surrogate posterior from the previous round is used to generate the training set of the next round. Here, the compute expended for simulating the next round training set can be directly used for calculating the importance weights for the current round NDE posterior. Thus, each round of SNPE contributes to the final effective sample size at no extra cost. The ratio between the effective and total sample size (sampling efficiency) for each round becomes an indicator for the convergence of the NDE posterior towards the exact posterior, and thus an early stopping criterion. Once the surrogate posterior stops improving for the next round, one may directly acquire more effective posterior samples simply by importance sampling from the previous round surrogate posterior with the highest sampling efficiency. The full SNPE with importance sampling integration (SNPE-IS) algorithm is presented in Algorithm 1.
@@
2:  Output: posterior samples $\{\theta_{1:N\cdot R}\}$, weights $\{w_{1:N\cdot R}\}$, effective sample sizes $\{n_{\rm eff,1:R}\}$, final surrogate posterior $\hat{q}(\theta)$
@@
14:     Train $\hat{q}(\theta)$ on $\{\theta_{1:N},x_{1:N}\}$ using Negative-Log-Likelihood (NLL) loss
@@ 4 Discussion
We consider the current version (v0.4.1) of nbi “production-ready” for inference problems involving sequential data, such as light curves and spectra. The only customization we require is a “noise” function that captures the instrumental effects of the underlying dataset. In a companion paper (Zhang et al., 2023), we apply nbi to the problem of stellar spectra fitting, and discuss how nbi facilitates easy the distribution of the trained amortized inference models. Looking forward, we plan to include more types of featurizer networks to allow for the out-of-the-box utility of nbi to be extended to more types of data. One aspect worth mentioning is specialized networks for irregularly sampled light curve. We point out that for single inference tasks utilizing SNPE, there is no distinction between regularly and irregularly sampled data, and the 1D ResNet-GRU network provided with nbi should be sufficient. We outline several aspects for an extended version of this work.